# Mamba-based Deep Learning Approaches for Sleep Staging on a Wireless Multimodal Wearable System without Electroencephalography

### v1

@@ 2.4 Training
The training, validation, and testing sets were randomly split in a size ratio to sizes . During the training process, recordings were clipped to random segments to prevent the model from overfitting to the high-level structure of most recordings, which mostly consisted of wake periods at the start and end of the recording with sleep in the middle. Cross-entropy loss was weighted on inverse class frequency to adjust for the moderate class imbalance as encapsulated by Figure 1C. The model from the iteration that maximized balanced validation accuracy was chosen as the final model. Both CRNN and RNN models were trained with the Adam optimizer [32].
The CRNN model was trained with single batches. Recordings were clipped to segments of size varying uniformly between and , and age/sex values individually were set to with probability , an arbitrary value, as a form of data augmentation. The RNN model was trained on a batch size of , with recordings clipped randomly to sequence lengths of seconds.
@@ 2.5 Ensembling
As decision tree-based architectures do not output class-wise logits, epoch-wise predictions on the RNN model were generated by aggregating 30 one-second predictions and reducing with the mode.
@@ 3 Results
Results are compiled of 3-class, 4-class, and 5-class sleep staging for the RNN and CRNN architectures and for regular and ensemble models. Overall performance is evaluated by aggregating all test set recordings and computing the balanced accuracy (B.Acc.), precision (Prec.), recall (Rec.), F1, Cohen’s Kappa (), and Matthews correlation coefficient (MCC). We additionally conduct analyses of averaged per-recording metrics, biases in the amount of each sleep stage, and differences in performance across age, PLMI, AHI, BMI, TST, and sleep efficiency. Because individual recordings do not necessarily contain instances of all five sleep stages, balanced accuracy may not be a sensible metric for recording-wise comparisons; hence, accuracy (Acc.) is presented in lieu of balanced accuracy for those cases.
@@ 3.1 Model Evaluation
Figure 5 displays the confusion matrices for all -class ensemble models and overall metrics across all -class regular and ensemble models. No significant deviations in performance were observed in the RNN models as tested on 1-second windows and as tested on 30-second windows, of which the latter is presented in this paper. The RNN ensemble model yields the best metrics and dominates most class-wise metrics (Appendix D). Ensemble models perform better than their single-model counterparts, while RNNs perform better than the CRNNs despite the shallower feature set and smaller parameter count. Most misclassifications are made between N1 and every other class, as well as between N2 and N3 for 5-class sleep staging.
Lastly, in a set of young and healthy sleepers (age < 40, AHI < 5, PLMI < 5), the RNN ensemble model’s performance is significantly boosted as denoted in the confusion matrices and metrics of Figure 6.
![[Uncaptioned image]](extracted/6084238/figures/ensemble_outputs.png)
@@
![[Uncaptioned image]](extracted/6084238/figures/healthy-subset.png)
Metric (%) Classes Balanced Accuracy Precision Recall F1 Cohen’s MCC 5 69.26 76.85 70.07 71.55 60.48 61.60 4 79.39 78.82 77.00 76.89 67.23 67.92 3 86.32 87.63 86.70 86.88 77.87 78.14
@@ 3.2 Global Sleep Metrics
As a means of evaluating model bias and verifying model robustness, Figure 7 lays out scatter plots and Bland-Altman plots for the 5-class RNN ensemble, comparing RNN predicted vs. PSG actual values for a number of summary metrics including time asleep, % N1 sleep,% N2 sleep, % N3 sleep, and % REM sleep. The model trends towards slightly underestimating N2 sleep and slightly overestimating N1 sleep and wake.
@@ 3.3 Performance by Clinical Group
Averaged recording-wise metrics are cataloged in Table 2 for the 5-class ensemble CRNN and RNN, wherein linear regressions were performed to fit various recording metrics to accuracy. Performance of the RNN ensemble is robust to differences in sex, periodic limb movement index, body mass index, total sleep time and sleep efficiency, but shows slightly worse performance with greater age and higher AHI. The CRNN model is robust to differences in age, sex, PLMI, and BMI but shows slightly worse performance with higher AHI, shorter TST, and lower sleep efficiency. Notably, even where performance varies by clinical factors, the percentage of explained change in performance is nevertheless weak for all metrics (). There is no significant difference in accuracy between the raters in the testing set (RNN , CRNN , one-way F-test).
| RNN | CRNN | ||||||
| Metric (%) | |||||||
| Accuracy | 65.09 | 9.84 | 62.58 | 12.50 | |||
| F1 | 67.69 | 9.92 | 63.81 | 14.17 | |||
| Cohen’s | 50.39 | 14.91 | 46.33 | 16.13 | |||
| Matthews Correlation Coefficient | 52.61 | 13.94 | 48.77 | 15.91 | |||
| Sleep Statistic | |||||||
| Age | 0.0827 | 0.0350 | 0.0346 | 0.1779 | |||
| Sex | 0.0689 | 0.0552 | 0.024 | 0.2630 | |||
| Periodic Limb Movement Index (PLMI) | 0.0120 | 0.4300 | 0.0012 | 0.8061 | |||
| Apnea-Hypopnea Index (AHI) | 0.1525 | 0.0350 | 0.1622 | 0.0025 | |||
| Body Mass Index (BMI) | 0.0164 | 0.3559 | 0.0000 | 0.9824 | |||
| Total Sleep Time (TST) | 0.0000 | 0.9921 | 0.0853 | 0.0322 | |||
| Sleep Efficiency | 0.0045 | 0.6316 | 0.1503 | 0.0038 | |||
@@ 4.1 Key Findings
This work has successfully developed Mamba-based CRNN and RNN models to accurately perform inference from the ANNE One device for major sleep stages. The models demonstrated good performance across a wide range of patient characteristics, without a need to screen and exclude recordings and epochs on the basis of signal quality, and with the RNN architecture slightly outperforming the CRNN architecture.
@@ 4.2 Related Studies
Previous studies have undertaken the challenge of inferring sleep stage from non-EEG signals (Tables 3 & 4) using a mixture of cardiopulmonary (ECG, PPG, respiratory impedance bands, respiratory flow cannulas) and accelerometry data. Many studies used data from conventional PSG recordings [20, 21, 22, 23, 24, 25, 26] although a handful of studies used data from purpose-designed wearable sensors [28, 21, 19, 23, 24, 27]. In general, models trained on and applied to PSG data performed better than models utilizing wearable sensor data. This is not surprising since signal quality is often much higher for attended in-laboratory PSGs where more intrusive sensors can be used, and signal quality is continuously monitored and issues can be addressed by the technologist as they occur. Even so, despite using wearable rather than PSG sensor data, the performance of our ensemble RNN model approached [21, 24, 25, 26, 20, 22] or exceeded [23] that of models trained on much larger sets of PSG data.
Our model exceeded the overall performance of most wearable sensor based models [27, 23, 28] with the exception of one [24]. Of note, this study examined a younger set of participants (mean age 45) who were free of psychiatric and neurological co-morbidity. In contrast, we studied a much older (mean age 57.49) participants with significant sleep co-morbidity (45.81% with AHI > 5). When we limit our analyses to younger individuals without sleep apnea or significant periodic limb movements, model performance improves further (Figure 6). Another contrast between [24] and the present study is that [24] performed much stricter quality control for recording quality with 54/136 or nearly 40% of nights of wearable sensor recording excluded from analysis due to poor signal quality. In contrast, in the interest of evaluating real-world performance, we did not exclude any recording or any epochs on the basis of signal quality and the reported results reflect consideration of all available epochs of recording, irrespective of signal quality. The robustness of our model to missing data or poor signal quality may be related in part to its use of multiple streams of data such that sensors returning poor quality data are compensated for in the model by those returning good quality data.
Of particular comparative interest is the one previous paper addressing sleep stage inference from ANNE One data [19]. Ultimately, success in developing a population model was limited, and while there was some success in generating person-specific models, the performance of our ensemble RNN model exceeded even the person-specific models generated in this paper.
@@ 4.3 Model Performance
Both model architectures obtain maximal per-class performances for wake and REM, and minimal per-class performances for N1 and N2 sleep. As a transitory state, periods of N1 sleep are short and instable, making its discrimination a difficult task. Physiological similarities between N2 and N3 sleep, such as cardiorespiratory signatures, similarly complicate the model’s ability to differentiate them, as evidenced by N2-N3 confusion as the most frequent misprediction for 5-class models (as well as N1/N2-N3 confusion for 4-class models). Conversely, wake and REM sleep have clearer differentiating characteristics, such as higher heart rate, which allows for easier classification.
@@
No statistically significant variations in performance based on PLMI, BMI, TST, or sleep efficiency were exhibited by the RNN 5-class ensemble model, providing an important advantage when applied in clinical contexts. The significant () association discovered between AHI and accuracy may owe to the fact that individuals with sleep apnea experience sleep fragmentation, resulting in more time in N1 sleep, one of the most-difficult stages to separate; even so, the association was modest (). Model performance also correlates between age and accuracy (), though this may in part be explained by the weak agreement between AHI and age ().
@@ 4.4 Limitations, Strengths, and Future Directions
A few methodological limitations are worth considering. 1) The differing nature of the way our CRNN and RNN models handle preprocessing and predictions (notably, the 30-second window employed by the CRNN as opposed to the 1-second window used in the RNN) make the models difficult to compare; our reconciling of the RNN to make predictions for 30-second windows may not reflect its full performance. 2) As neural networks, CRNNs and RNNs provide limited insight into the decoding of sleep stages from the given sensor data. Further analyses targeted toward interpreting the neural networks may be explored. 3) The recordings were overnight with a brief period of wakefulness before lights out, resulting in a preponderance of sleep vs. wake. It is difficult to know for sure how the model will perform in detecting daytime naps during long ambulatory recordings and further studies with ambulatory recordings in clinical populations will be informative.
This study also had many strengths: 1) The ANNE One system itself, with multiple channels of data recorded from a pair or minimally intrusive sensors, is a strength insofar as it provides an element of redundancy should one sensor return low quality signal, 2) the combination of Mamba with an RNN architecture allowed the achievement of relatively high performance with a relatively parameter-scarce architecture, and 3) the study population was varied, including participants from across a spectrum of sleep disorders.
@@ Acknowledgments
Appendix A: Bidirectional Mamba Block
Appendix B: Scalar and Frequency Features
@@
Appendix D: Class-Wise Evaluation Metrics
5-Class Sleep Staging 4-Class Sleep Staging 3-Class Sleep Staging B.Acc. Prec. Rec. F1 MCC B.Acc. Prec. Rec. F1 MCC B.Acc. Prec. Rec. F1 MCC Wake Wake Wake CRNN 83.52 84.10 84.13 84.07 67.46 67.53 84.59 84.80 84.75 84.77 69.01 69.02 84.57 84.70 84.45 84.50 68.57 68.70 RNN 86.00 88.85 87.59 87.29 74.01 75.63 88.00 89.57 89.08 88.94 77.36 78.05 88.67 89.26 89.23 89.18 77.90 78.01 CRNN Ens. 85.63 86.90 86.64 86.49 72.36 72.83 86.96 87.40 87.41 87.38 74.24 74.28 88.05 88.52 88.52 88.49 76.49 76.55 RNN Ens. 86.68 88.63 87.95 87.75 74.93 75.88 88.49 89.14 89.10 89.04 77.61 77.74 89.03 89.41 89.42 89.39 78.36 78.39 N1 Sleep N1/N2 Sleep Non-REM Sleep CRNN 66.02 86.62 79.98 82.65 22.47 24.23 62.09 67.98 68.51 65.69 26.58 29.20 79.70 80.20 79.98 79.88 59.68 60.00 RNN 67.77 87.06 77.97 81.43 22.61 25.40 68.42 71.65 72.13 71.34 38.39 39.00 81.68 82.36 82.00 81.89 63.72 64.18 CRNN Ens. 68.82 87.42 80.47 83.18 25.87 28.12 71.45 74.71 75.07 74.32 44.79 45.56 82.29 82.67 82.52 82.46 64.83 65.05 RNN Ens. 68.96 87.42 79.35 82.44 24.96 27.61 73.52 76.06 76.40 75.96 48.45 48.83 85.15 85.35 85.30 85.28 70.47 70.56 N2 Sleep CRNN 65.60 74.43 76.10 74.58 34.46 35.48 RNN 64.82 76.06 77.46 74.93 34.63 37.49 CRNN Ens. 65.29 75.54 77.14 75.00 35.03 37.10 RNN Ens. 68.12 77.33 78.59 77.02 40.56 42.13 N3 Sleep N3 Sleep CRNN 73.95 90.34 89.41 89.83 44.03 44.26 75.29 89.62 81.09 84.14 31.74 35.77 RNN 82.25 91.85 87.58 89.08 47.45 50.21 82.22 91.58 84.59 86.94 42.11 46.62 CRNN Ens. 77.66 91.19 89.59 90.25 48.04 48.69 78.35 91.05 88.33 89.39 45.80 47.15 RNN Ens. 81.20 92.14 90.30 91.01 52.80 53.78 82.18 92.12 89.32 90.35 51.18 52.91 REM Sleep REM Sleep REM Sleep CRNN 72.17 91.05 91.33 91.18 45.86 45.90 73.16 90.10 87.77 88.75 38.54 39.38 78.98 92.24 91.29 91.70 52.43 52.79 RNN 82.42 93.23 92.29 92.67 58.19 58.66 82.85 93.17 91.90 92.40 57.33 58.04 87.46 93.49 88.87 90.37 52.70 56.55 CRNN Ens. 81.70 92.69 91.06 91.70 54.01 54.92 83.06 92.82 90.54 91.38 53.64 55.13 83.47 92.73 89.69 90.78 51.80 53.90 RNN Ens. 84.79 94.20 93.63 93.86 64.41 64.68 87.16 94.40 93.22 93.65 64.47 65.33 88.17 94.49 92.94 93.48 61.16 65.42

### v2

@@ 2.4 Training
The training, validation, and testing sets were randomly split in a ratio to sizes . During the training process, recordings were clipped to random segments to prevent the model from overfitting to the high-level structure of most recordings, i.e., wake periods at the start and end with sleep in the middle. Cross-entropy loss was used, weighted on inverse class frequency to adjust for the moderate class imbalance as encapsulated by Figure 1C. The model from the training epoch that maximized balanced validation accuracy was chosen as the final model. Both CRNN and RNN models were trained with Adam [32], a widely-used adaptive learning rate optimizer.
The CRNN model was trained on single batches. Recordings were clipped to segments of size varying uniformly between and , and age/sex values individually were set to with probability , an arbitrary value, as a form of data augmentation. The RNN model was trained on a batch size of , with recordings clipped randomly to sequence lengths of seconds.
@@ 2.5 Ensembling
2.6 Statistical Analyses
Results were compiled of 3-class, 4-class, and 5-class sleep staging for the RNN and CRNN architectures and for regular and ensemble models. Overall performance was evaluated by aggregating all test set recordings and computing the balanced accuracy, weighted precision, weighted recall, weighted F1, Cohen’s Kappa (), and Matthews correlation coefficient (MCC).
For the best model type, performance was plotted on a healthy subset of the test set and ablation studies are carried out using various subsets of the ANNE One sensor data and a different sequential architecture. For the best 5-class model, Blant-Altman plots were created to evaluate class-wise biases, followed by analyses of averaged per-recording metrics both individually and across clinical variables including age, sex, PLMI, AHI, BMI, TST, and sleep efficiency. Because individual recordings do not necessarily contain instances of all five sleep stages, balanced accuracy may not be a sensible metric for recording-wise comparisons; hence, accuracy was presented in lieu of balanced accuracy for those cases.
@@ 3.1 Model Evaluation
Figure 5 displays the confusion matrices for all -class ensemble models and overall metrics across all -class regular and ensemble models. No meaningful deviations in performance were observed in the RNN models as tested on 1-second windows and as tested on 30-second windows (balanced accuracies of 1-second RNN ensembles are 83.74%, 75.06%, 64.86% for 3/4/5-class models), of which the latter is presented in this paper. The RNN ensemble model yields the best metrics and dominates most class-wise metrics (5-class balanced accuracies of 86.19%, 69.62%, 67.72%, 81.16%, and 86.60% for wake, N1, N2, N3, and REM sleep respectively, Appendix F). Ensemble models perform better than their single-model counterparts, while RNNs perform better than the CRNNs despite the shallower feature set and smaller parameter count. Most misclassifications are made between N1 and every other class, as well as between N2 and N3 for 5-class sleep staging.
@@
In a set of young and healthy sleepers (age < 40, AHI < 5, PLMI < 5), the RNN ensemble model experiences a significant boost in performance as denoted in Figure 6 although removal of these participants from the test set does not substantially worsen model performance, indicating that overall model performance is not driven by these healthy participants (Appendix G, Figure 12).
Metric (%) Classes Balanced Accuracy Precision Recall F1 Cohen’s MCC 5 69.41 75.30 68.73 69.58 59.83 61.17 4 79.89 78.40 75.83 75.57 66.42 67.58 3 87.67 88.18 87.69 87.78 79.09 79.29
3.2 Feature Ablation Study
We next investigated which sets of sensors (chest module vs. limb module; ECG vs. PPG. vs. other) were most influential to model performance. To do so, modified instances of the RNN ensemble model were trained to determine the utility of its components. Firstly, feature sets solely from the ANNE One’s chest module (ECG, accelerometry, chest temperature) or limb module (PPG, limb temperature) were examined. As shown in Table 2, the chest-only model trails closely behind the original in performance while the limb-only model exhibits a marked degradation () for all metrics. This strongly suggests that the chest module is the principal driver behind model accuracy, although the limb module alone is capable of recovering sleep staging information to a rough degree. Next, because the ECG and and PPG sensors themselves are the most likely to experience poor signal quality, we examined the performance of RNN ensemble models trained on only the accelerometry and temperature sensors. Remarkably, exclusion of ECG and PPG data resulted in minimal loss of accuracy, meaning that information in the ECG and PPG is also being captured effectively by the accelerometry signals.
[…]
3.3 Recurrent Architectures
In order to understand whether the incorporation of Mamba improved the RNN ensemble model, a similarly-sized variant model replacing Mamba with LSTM was trained, its results provided by Table 3. The LSTM model shows a minor loss in all metrics, confirming the Mamba approach is beneficial.
[…]
@@ 3.4 Global Sleep Metrics
As a means of evaluating model bias and verifying model robustness, Figure 7 lays out scatter plots and Bland-Altman plots for the 5-class RNN ensemble, comparing RNN predicted vs. PSG actual values for a number of summary metrics including time asleep, % N1 sleep, % N2 sleep, % N3 sleep, and % REM sleep. The model trends towards slightly underestimating wake and N2 sleep and slightly overestimating N1 and N3 sleep.
@@ 3.5 Performance by Clinical Group
Averaged recording-wise metrics are catalogued in Table 4 for the 5-class RNN ensemble, wherein linear regressions were performed to fit various recording metrics to accuracy (plots in Appendix G, Figure 13). Performance of the RNN ensemble is robust to differences in all metrics except AHI, which had a weak negative correlation with accuracy, though the proportion of variance in accuracy explained by AHI is modest (RNN ). Four annotators were involved in generating the ground truth labels used in this study and the agreement between the RNN model and the ground truth labels did not differ by the individual who generated the annotations (RNN , one-way F-test).
Metric (%) | ||||
Accuracy | 64.54 | 10.77 | ||
F1 | 66.72 | 10.28 | ||
Cohen’s | 50.58 | 14.50 | ||
Matthews Correlation Coefficient | 52.92 | 13.40 | ||
Sleep Statistic | ||||
Age | 0.0618 | 0.0700 | ||
Sex | 0.0334 | 0.1861 | ||
Periodic Limb Movement Index (PLMI) | 0.0338 | 0.1830 | ||
Apnea-Hypopnea Index (AHI) | 0.0751 | 0.0449 | ||
Body Mass Index (BMI) | 0.0350 | 0.1753 | ||
Total Sleep Time (TST) | 0.0055 | 0.5940 | ||
Sleep Efficiency | 0.0002 | 0.9281 | ||
@@ 4.1 Key Findings
This work has developed Mamba-based CRNN and RNN models to accurately perform inference from the ANNE One device for major sleep stages. The models demonstrate good performance across a wide range of patient characteristics, without a need to screen and exclude recordings and epochs on the basis of signal quality, and with the RNN architecture slightly outperforming the CRNN architecture.
@@ 4.2 Related Studies
Previous studies have undertaken the challenge of inferring sleep stage from non-EEG signals (Tables 5 & 7) using a mixture of cardiopulmonary (ECG, PPG, respiratory impedance bands, respiratory flow cannulas) and accelerometry data. Many studies used data from conventional PSG recordings [20, 21, 22, 23, 24, 25, 26] although a handful of studies used data from purpose-designed wearable sensors [28, 21, 19, 23, 24, 27]. In general, models trained on and applied to PSG data performed better than models utilizing wearable sensor data. This is not surprising since signal quality is often much higher for attended in-laboratory PSGs where more intrusive sensors can be used, and signal quality is continuously monitored and issues can be addressed by the technologist as they occur. Even so, despite using wearable rather than PSG sensor data, the performance of our ensemble RNN model approached [21, 24, 25, 26, 20, 22] or exceeded [23] that of models trained on much larger sets of PSG data.
Our model exceeded the overall performance of most wearable sensor based models [27, 23, 28] with the exception of one [24]. Of note, this study examined a younger set of participants (mean age 45) who were free of psychiatric and neurological co-morbidity, which may influence the relationship between sleep stage and its autonomic, cardiac, and pulmonary manifestations. In contrast, we studied an older (mean age 57.49) set of participants with significant sleep co-morbidity (45.81% with AHI > 5). When we limit our analyses to younger individuals without sleep apnea or significant periodic limb movements, model performance improves further (Figure 6). Another contrast between [24] and the present study is that [24] performed much stricter quality control for recording quality with 54/136 or nearly 40% of nights of wearable sensor recording excluded from analysis due to poor signal quality. In contrast, in the interest of evaluating real-world performance, we did not exclude any recording or any epochs on the basis of signal quality and the reported results reflect consideration of all available epochs of recording, irrespective of signal quality. The robustness of our model to missing data or poor signal quality is likely related in part to its use of multiple data streams, in particular accelerometry, which appeared to capture much of the information that would otherwise have been captured by ECG or PPG.
Of particular comparative interest is the one previous paper addressing sleep stage inference from ANNE One data [19]. Ultimately, success in developing a population model was limited, and while there was some success in generating person-specific models, the performance of our ensemble RNN model exceeded even the best person-specific model in the paper which had a 4-class balanced accuracy of 73.4%.
@@ 4.3 Model Performance
Both CRNN and RNN model architectures obtain maximal per-class performances for wake and REM, and minimal per-class performances for N1 and N2 sleep. As a transitory state, periods of N1 sleep are short and instable, making its discrimination a difficult task. Physiological similarities between N2 and N3 sleep, such as cardiorespiratory signatures, similarly complicate the model’s ability to differentiate them, as evidenced by N2-N3 confusion as the most frequent misprediction for 5-class models (as well as N1/N2-N3 confusion for 4-class models). Ultimately, the differences betwen N1, N2, and N3 represent a matter of depth rather than kind, and some degree of difficulty in distinguishing these without EEG is to be expected. Conversely, wake and REM sleep have clearer differentiating characteristics, such as higher heart rate, which allows for easier classification.
@@
No statistically significant variations in accuracy based on age, sex, PLMI, BMI, TST, or sleep efficiency were exhibited by the RNN 5-class ensemble model, providing an important advantage when applied in clinical contexts. The marginally significant () association between AHI and accuracy may owe to the fact that individuals with sleep apnea experience sleep fragmentation, resulting in N1 sleep, one of the most difficult stages to separate; nonetheless, the association was modest ().
The series of ablation analyses establish that much of the RNN model’s performance was likely driven by chest-module features (i.e. ECG, temperature, and xyz accelerometry) rather than limb module (i.e. PPG and temperature), meaning that omitting the limb module may be reasonable if participant burden is a concern and collection of oximetry data is not necessary. Remarkably, exclusion of both ECG and PPG data with retention of only accelerometry and temperature data resulted in model performance almost indistinguishable from use of all the chest module sensors, suggesting that the inclusion of traditional heart rate sensors (PPG, ECG) is not necessary. This highlights the richness of accelerometry data obtained at the chest, spectrograms of which may capture both cardiac (ballistocardiogram) and respiratory data, as well as measures of movement. It may also contribute to the robustness of our model given the relative reliability of the accelerometers compared to the ECG/PPG sensors.
@@ 4.4 Limitations, Strengths, and Future Directions
This study had many strengths: 1) The ANNE One system itself, with multiple channels of data recorded from a pair or minimally intrusive sensors, is a strength insofar as it provides an element of redundancy should one sensor return low quality signal, 2) the inclusion of Mamba allowed the achievement of relatively high performance with a relatively parameter-scarce architecture, and 3) the study population was varied, including participants from across a spectrum of sleep disorders.
A few methodological limitations are also worth considering. 1) The differing nature of the CRNN and RNN in preprocessing and predicting (notably, the CRNN’s 30-second window as opposed to the RNN’s 1-second window) make the models difficult to compare; our reconciling of the RNN to make predictions for 30-second windows may not reflect its full performance. 2) The black box nature of CRNNs and RNNs make it difficult to ascertain the specific mechanisms used to infer sleep stage from the sensor data. Further analyses targeted toward interpreting the neural networks may be explored. 3) The recordings were overnight with a brief period of wakefulness before lights out, resulting in a preponderance of sleep vs. wake. It is difficult to know for sure how the model will perform in detecting daytime naps during long ambulatory recordings and further studies with ambulatory recordings in clinical populations will be informative. 4) The recordings were all obtained from a single site. Additional work is needed to assess the extent to which these results will be generalizable to ANNE recordings made in different settings on different patient populations.
@@ Acknowledgments
Appendix A: Summary Statistics by Groups
Variable Age 65 Age 65 p AHI AHI 5 p Age [y] 64.01 ± 14.34 51.82 ± 18.07 1.48 Sex [% Female] 44.52 ± 49.70 52.83 ± 49.92 0.123 40.36 ± 49.06 57.29 ± 49.47 1.34 AHI 14.23 ± 16.64 9.90 ± 17.59 0.0203 TST [h] 4.61 ± 1.29 5.43 ± 1.48 1.46 4.86 ± 1.35 5.30 ± 1.52 4.82 PLMI 23.34 ± 35.41 9.22 ± 16.82 7.96 16.64 ± 25.38 13.55 ± 28.18 0.281 Body Mass Index 27.89 ± 4.59 27.90 ± 6.19 0.986 29.45 ± 5.72 26.55 ± 5.13 7.54 Sleep Efficiency [%] 65.81 ± 16.40 75.84 ± 17.81 1.27 69.61 ± 17.48 73.61 ± 18.13 0.0356 Awake [%] 46.64 ± 15.00 37.10 ± 16.47 5.22 43.01 ± 16.08 39.25 ± 16.78 0.0321 N1 Sleep [%] 12.13 ± 6.64 10.43 ± 5.83 0.0110 13.15 ± 6.97 9.37 ± 4.88 5.41 N2 Sleep [%] 25.87 ± 10.31 30.06 ± 10.46 2.17 26.77 ± 11.03 29.72 ± 10.02 8.59 N3 Sleep [%] 8.20 ± 6.73 12.64 ± 8.48 2.38 9.00 ± 7.43 12.41 ± 8.34 6.36 REM Sleep [%] 7.16 ± 5.20 9.77 ± 6.40 5.96 8.07 ± 5.84 9.25 ± 6.22 0.0679
Appendix B: Bidirectional Mamba Block
@@
Appendix D: Scalar and Frequency Features
@@
Appendix F: Class-Wise Evaluation Metrics
5-Class Sleep Staging 4-Class Sleep Staging 3-Class Sleep Staging Regular Ensemble Regular Ensemble Regular Ensemble Metric (%) CRNN RNN CRNN RNN CRNN RNN CRNN RNN CRNN RNN CRNN RNN Wake Wake Wake Balanced Accuracy 83.75 84.93 86.47 86.19 84.13 86.12 87.93 87.87 78.57 88.14 87.97 88.27 Precision 84.61 88.16 87.64 88.44 84.48 88.57 88.34 88.67 80.48 88.70 88.24 88.82 Recall 84.68 87.28 87.64 88.04 83.38 88.07 88.32 88.70 76.26 88.72 88.12 88.85 F1 84.63 86.93 87.54 87.81 83.53 87.82 88.33 88.66 76.40 88.71 88.16 88.83 Cohen’s 67.92 72.53 73.93 74.41 66.37 74.42 75.74 76.32 53.63 76.47 75.45 72.89 MCC 67.95 73.94 74.11 75.17 67.04 75.32 75.74 76.37 56.43 76.48 75.49 76.74 N1 Sleep N1/N2 Sleep Non-REM Sleep Balanced Accuracy 62.71 65.07 67.99 69.62 69.22 67.11 71.58 73.47 72.35 83.46 81.03 85.42 Precision 85.30 85.70 86.74 87.09 73.32 68.97 75.37 75.61 74.26 83.60 81.90 85.50 Recall 82.64 78.13 81.74 79.31 73.05 69.43 75.06 75.81 72.28 83.45 80.99 85.41 F1 83.84 81.14 83.77 82.25 71.73 69.01 74.00 75.36 71.73 83.43 80.87 85.41 Cohen’s 21.74 20.73 27.21 26.46 40.67 34.97 45.39 48.25 44.64 66.90 62.01 70.83 MCC 22.12 22.58 28.61 29.06 42.52 35.19 47.03 48.74 46.54 67.05 62.90 70.92 N2 Sleep Balanced Accuracy 64.85 67.47 65.88 67.72 Precision 73.86 74.95 76.25 76.33 Recall 75.49 76.27 77.22 77.50 F1 73.44 75.03 74.80 75.82 Cohen’s 33.51 37.98 36.71 39.58 MCC 35.28 38.87 39.68 41.29 N3 Sleep N3 Sleep Balanced Accuracy 76.48 78.15 79.40 81.16 77.36 81.50 80.08 82.18 Precision 89.80 90.53 90.86 91.71 90.46 90.98 90.92 91.72 Recall 85.36 87.47 87.60 89.66 88.07 81.92 86.85 88.59 F1 87.03 88.64 88.82 90.45 89.02 84.87 88.32 89.72 Cohen’s 39.73 45.03 46.36 52.35 45.54 38.84 45.47 50.70 MCC 41.84 46.52 48.08 53.44 46.58 44.30 47.77 52.60 REM Sleep REM Sleep REM Sleep Balanced Accuracy 75.29 81.55 82.31 86.60 77.83 77.66 82.99 88.67 70.09 85.57 84.13 89.26 Precision 90.95 93.38 92.70 94.69 91.67 93.31 92.63 94.97 90.33 93.47 92.52 94.87 Recall 90.33 93.24 91.39 94.31 90.93 93.76 90.70 94.26 91.04 91.80 89.04 93.81 F1 90.61 93.31 91.91 94.46 91.26 93.46 91.42 94.52 90.63 92.40 90.25 94.18 Cohen’s 47.59 61.83 56.59 69.06 51.59 60.63 55.18 70.06 43.76 60.09 51.95 68.78 MCC 47.72 61.85 57.27 69.23 51.80 61.06 56.35 70.54 44.03 61.23 54.38 69.61
Appendix G: Additional Model Performance Plots
Metric (%) Classes Balanced Accuracy Precision Recall F1 Cohen’s MCC 5 63.56 70.91 63.34 65.30 51.43 52.60 4 73.68 75.40 73.15 73.73 60.11 60.46 3 83.13 84.02 83.13 83.38 71.28 71.41